Ned's New Wave Barber Shop specializes in modern unisex haircuts. The only service available at Ned's is a “20-minute” haircut for which the customer is charged $10. The shop has five (5) barbers. (Ned does not work in the shop and, as owner/entrepreneur, he takes no salary.) Each barber is paid an annual salary of $18,000. All equipment including store fixtures and barbering equipment is leased on an annual basis at $4,500 per year. Building space is leased at the rate of $500 per month (or $6,000 per year). Ned is concerned about the shop’s cost structure and seeks your advice.

Ned is considering the following changes to how he pays the barbers and the landlord.

a.   Change only how barbers are paid. Instead of receiving a salary, barbers would receive a commission equal to 50% of the selling price of each haircut. What is the new contribution margin per haircut? What is the new annual break-even point in number of haircuts?

b. Change only the building lease agreement so that the landlord receives monthly rent of $100 plus 10% of the revenue per haircut. What is the new contribution margin per haircut? What is the new annual break-even point in number ofhaircuts?

c. Change both the compensation system for the barbers and the building lease agreement. If both are changed, what is the new contribution margin per haircut? What is the new annual break-even point in number of haircuts?

Evaluate the four cost structures proposed above, i.e., the original cost structure and the three changes in question. Which cost structure would you recommend and why? Explain how your recommendation depends on your assumptions.

For example: a. Do you think Ned will be able to implement the changes outlined in question above?

Will the barbers agree to work on commission? Will the landlord agree to accept less fixed rent in exchange for a share of revenue?

b. Do you expect the number of haircuts sold to differ across the four cost structures? If so, what differences would you expect?

c. What are the differences in risk to Ned across the four cost structures?
Answer #1
a. Breakeven point = Fixed cost / Contributon margin per unit
$
Revenue per haircut 10
Commission to barbers (50% of revenue) 5
Contribution margin per haircut 5
Annual fixed cost $
Lease of equipment 4500
Lease of building 6000
10500
Breakeven point = 10500 / 5 = 2100 hair cuts
b.
$
Revenue per haircut 10
Lease of building (10% of revenue) 1
Contribution margin per haircut 9
Annual fixed cost $
Salary of barber (18000 *5) 90000
Lease of equipment 4500
Lease of building ($ 100 * 12 months) 1200
95700
Breakeven point = 95700 / 9 = 10634 hair cuts
c. $
Revenue per haircut 10
Commission to barbers (50% of revenue) 5
Lease of building (10% of revenue) 1
Contribution margin per haircut 4
Annual fixed cost $
Lease of equipment 4500
Lease of building ($ 100 * 12 months) 1200
5700
Breakeven point = 5700 / 4 = 1425 hair cuts
d. Original break even point
$
Revenue per haircut 10
Contribution margin per haircut 10
Annual fixed cost $
Salary of barber (18000 *5) 90000
Lease of equipment 4500
Lease of building 6000
100500
Breakeven point = 100500 / 10 = 10050 hair cuts
In my opinion option a is better i.e. giving commission to brabers and all other other expenses will be fixed.
Leaseing equipment company will not be ready to accept lease on the basis of performance of the barber shop.
Barbers will be ready to accept the commission on the basis of number of haircuts.
They will get commission of 50% of revenue i.e. $ 5 per hair cut. So they have to perform 18000 hair cuts to match their current remuneration.
So if number of haircuts are expected to be more than 18000 in a year, barbers will accept proposal 1.